OpenStudy (jet_black_heart1423):

The map of a walking trail is drawn on a coordinate grid with three points of interest. The trail starts at R(−1, 4) and goes to S(5, 4) and continues to T(5, −2). The total length of the walking trail is ____ units. (Input whole numbers only.)

OpenStudy (kittiwitti1):

Have you learned the Distance Formula? ☺

OpenStudy (jet_black_heart1423):

?

OpenStudy (kittiwitti1):

Has the Distance Formula been taught in class?

OpenStudy (jet_black_heart1423):

I'm homeschooled and my parents are at work

OpenStudy (kittiwitti1):

\[\text{Distance Formula: }y=\sqrt{\left(x_{2}-x_{1}\right)^{2}+\left(y_{2}-y_{1}\right)^{2}}\] Yes, but your course still gives -- Wait, you don't take K12?

OpenStudy (kittiwitti1):

I am under the impression that if it's an online school they're supposed to teach you these formulas. :v But anyway, the idea here is to use the Distance Formula that I just posted to find the lengths

OpenStudy (jet_black_heart1423):

no i have flvs

OpenStudy (kittiwitti1):

Hmm. I don't know what that is but I'll try my best to help ☺

OpenStudy (jet_black_heart1423):

and not all online schools teach that

OpenStudy (jet_black_heart1423):

okay please help

OpenStudy (kittiwitti1):



OpenStudy (jet_black_heart1423):

hey are you typing?

OpenStudy (kittiwitti1):

So your goal is to find the distance between each point, and then add those distances up so you get the full perimeter.  The Distance Formula uses two coordinates: (x1, y1) and (x2, y2) So if we were getting the distance of RS we would get (-1,4) and (5,4) Inputting this into the Distance Formula:\[Distance=\sqrt{\left(x_{2}-x_{1}\right)^{2}+\left(y_{2}-y_{1}\right)^{2}}\]Which then becomes:\[D=\sqrt{\left(5-(-1)\right)^{2}+\left(4-4\right)^{2}}\]